The 60-MHz 'H NMR spectrum of 1,2-diphenylethylami
Answer #1

1)Ha Proton(pink):

It will be most downfield peak after aromatic protons as it is a 30, benzylic, and has direct –ve inductive effect of –NH2 group. The C atom bearing this proton is stereogenic center. Chemical shift for this proton is 4.1 ppm. Splitting pattern for this proton is doublets of doublet (dd) (Its explanation given in point-2).

Splitting pattern for this proton with chemical shift in Hz are as follows,

ν1 = 235.2 Hz

ν2 = 241.2 Hz

ν3 = 245.4 Hz

ν4 = 250.8 Hz

ν2 - ν1 = 6Hz and ν4- ν3 = 5.4 Hz

i.e. coupling constant J = 5.8 H as average.

2)Hb(green) and Hc (Blue) :

These are diastereotopic protons i.e. replacement of any of these two protons by completely different group or atom will generate stereogenic center at C bearing them and this will lead to the formation of diastereomers hence they are termed as diastereotopic protons.

Diastereotopic protons have different chemical shifts i.e. they are chemically non-equivalent and hence they behave differently with Ha proton. This is reflected in there splitting pattern. Chemical shifts in Hz for these protons are as,

ν1 =148.5

ν2 = 155.5

ν3 =161.7

ν4 =166.2

ν5 = 168.7

ν6 = 171.6

ν7= 179.4

ν8 = 184.8

ν3 - ν1 =6.2 Hz (Vicinal coupling 1,3 coupling).

ν4- ν2 = 10.7 H(Geminal coupling).

ν8 - ν7 = 5.4 Hz.

Such a 3 kinds of coupling indicates doublets of doublets of triplet (ddt) some peaks with accidental overlap.

Hence the ABX pattern for given molecule.

====================================================================================